There were 104,830 people who attended a rock festival. if there were 8110 more boys than girls, and
24,810 fewer adults over 50 years of age than there were girls, how many of each group attended the festival.

Answer by JoelSchwartz(130)

x=number of boys who attended the rock festival
y=number of girls who attended the rock festival
z=number of adults over the age of fifty that attended the rock festival.
x=8,110+y
y-24,810=z
x+y+z=104,830
8,110+y+y+y-24,810=104,830
3y=121,530
y=40,510
x=48,620
z=15,700